# On computability of equilibrium states

### v2

@@ 2.2. Basic concepts in ergodic theory
where $\mathcal{V}$ ranges over all covers of $X$ contained (in the sense of inclusion) in $(\mathcal{V}_{n})_{T}^{m}$. In particular, if the potential $\varphi$ is identically zero, then the topological pressure $P(T,\varphi)$ is also called the topological entropy of $T$, and is denoted by $h_{\operatorname{top}}(T)$.
We refer to [PU10, Subsection 3.2] to demonstrate that the topological pressure is well-defined by (2.5). Let $(X,\rho)$ be a compact metric space. For all continuous map $T$ of $X$ and $T$-invariant Borel probability measure $\mu\in\mathcal{M}_{T}(X)$, one can define measure-theoretic entropy $h_{\mu}(T)$ of $T$ with respect to $\mu$ (see [PU10, Chapter 2.4]). In particular, assume that $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2, and $\mu$ is a $T$-invariant Borel probability measure. Recall that a measurable function $J\colon X\rightarrow[0,+\infty)$ is a Jacobian of $T$ with respect to $\mu$ if for each $\mu$-measurable set $A\subseteq X$ on which $T$ is injective, $T(A)$ is still $\mu$-measurable and $\mu(T(A))=\int_{A}\!J\,\mathrm{d}\mu.$ By [PU10, Proposition 2.9.5], there exists a unique Jacobian $J_{\mu}$ of $T$ with respect to $\mu$. Then the measure-theoretic entropy of $T$ with respect to $\mu$ has the following explicit expression $h_{\mu}(T)=\int\!\log(J_{\mu})\,\mathrm{d}\mu$ (see for example, [PU10, Proposition 2.9.7]). Under the above notations, one has the following theorem (see for example, [PU10, Theorem 3.4.1]).
@@
Let $(X,\rho)$ be a compact metric space. Then for each continuous map $T\colon X\rightarrow X$ and each potential $\varphi\in C(X)$, we have
$P(T,\varphi)=\sup\left\{h_{\mu}(T)+\int\!\varphi\,\mathrm{d}\mu:\mu\in\mathcal{M}_{T}(X)\right\}.$
In particular, if $\varphi\equiv 0$, then $h_{\operatorname{top}}(T)=\sup\{h_{\mu}(T):\mu\in\mathcal{M}_{T}(X)\}$.
@@
Let $(X,\rho)$ be a compact metric space. Then for each continuous map $T\colon X\rightarrow X$ and each potential $\varphi\in C(X)$, we say that a measure $\mu\in\mathcal{M}_{T}(X)$ is an equilibrium state of the map $T$ for the potential $\varphi$ if
@@
In particular, for $\varphi\equiv 0$, an equilibrium state of $T$ for $\varphi$ is also called a measure of maximal entropy of $T$.
@@ 2.3. Ruelle–Perron–Frobenius operators and Gibbs states
In this subsection, we review some notions and results for the Ruelle–Perron–Frobenius operator, the Gibbs state, and the equilibrium state. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ is a continuous function. Recall that the Ruelle operator $\mathcal{L}_{\varphi}$ acting on $C(X)$ is given by
@@
for each $u\in C(X)$ and each $x\in X$. By [PU10, Theorem 5.2.8], there exists an eigenmeasure $m\in\mathcal{P}(X)$ and an eigenvalue $c>0$ of the adjoint operator $\mathcal{L}_{\varphi}^{*}$ of $\mathcal{L}_{\varphi}$, i.e., $\mathcal{L}^{*}_{\varphi}(m)=cm$. Moreover, by [PU10, Proposition 5.2.11], the constant $c$ is precisely $\exp(P(T,\varphi)).$ It is convenient to consider the normalized Ruelle–Perron–Frobenius operator $\mathcal{L}_{\overline{\varphi}}$ with $\overline{\varphi}\coloneqq\varphi-P(T,\varphi).$ By (2.6), we have
@@
(2.8) $C^{-1}\leqslant\frac{\mu(T^{-n}_{x}(B(T^{n}(x),\xi)))}{\exp(S_{n}\varphi(x)-Pn)}\leqslant C$
for all $x\in X$ and $n\in\mathbb{N}.$ The following result (see for example, [PU10, Propositions 4.4.3, 5.1.5, 5.2.10, Corollaries 5.2.13, Lemma 5.6.1, and Theorem 5.6.2]) implies that a $T$-invariant Gibbs state is an ergodic equilibrium state of $T$ for the potential $\varphi$.
@@
Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2. Then for all $\mu\in\mathcal{M}_{T}(X)$ and Hölder continuous potential $\varphi\colon X\rightarrow\mathbb{R}$, the following statements are equivalent:
@@
    $\mu$ is an ergodic equilibrium state of $T$ for $\varphi$.
@@
The existence and uniqueness of the subsequential limit $u_{\varphi}$ is well-known. We include a proof in Corollary 2.9 together with some quantitative bounds specialized for the purpose of Computable Analysis. In the remainder of this subsection, we quantitatively re-develop some familiar facts about the Ruelle–Perron–Frobenius operators from the perspective of Computable Analysis in preparation for more technical analysis in Section 4.
Definition 2.6.
Let $(X,\rho)$ be a compact metric space, and $v_{0}\in(0,1]$ and $a_{0}>0$ be constants. A function $\varphi\colon X\rightarrow\mathbb{R}$ is Hölder continuous (with respect to the metric $\rho$) with exponent $v_{0}$ and constant $a_{0}$ if $\lvert\varphi(x)-\varphi(y)\rvert\leqslant a_{0}\rho(x,y)^{v_{0}}$ for all $x,\,y\in X$.
@@
Fix arbitrary constants $\eta,\,\xi,\,a_{0},\,v_{0}>0$, and $\lambda>1$. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2 with constants $\eta$, $\lambda$, and $\xi$, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ be Hölder continuous with respect to the metric $\rho$ with exponent $v_{0}$ and constant $a_{0}$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$, and $a\coloneqq\frac{a_{0}}{\lambda^{v_{0}}-1}$. Then
(2.9) $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\leqslant\exp(a\rho(x,y)^{v_{0}})\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)$
for each $n\in\mathbb{N}$ and each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$.
@@
Since $T$ is an open map that is distance-expanding with constants $\eta$, $\lambda$, and $\xi$. Then for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$ and each $\overline{x}\in T^{-n}(x)$, one can find a point $y(\overline{x})\coloneqq T_{\overline{x}}^{-n}(y)\in T^{-n}(y)$ satisfying that $\rho(T^{i}(\overline{x}),T^{i}(y(\overline{x})))\leqslant\lambda^{i-n}\rho(x,y)\leqslant\xi$ for each integer $10\leqslant i\leqslant n-$ 1. Noting that $\overline{\varphi}(x)=\varphi(x)-P(T,\varphi)$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$, we have
$S_{n}\overline{\varphi}(\overline{x})-S_{n}\overline{\varphi}(y(\overline{x}))$ $=\sum_{i=0}^{n-1}(\overline{\varphi}(T^{i}(\overline{x}))-\overline{\varphi}(T^{i}(y(\overline{x}))))$
$\leqslant\sum_{i=0}^{n-1}a_{0}\rho(T^{i}(\overline{x}),T^{i}(y(\overline{x})))^{v_{0}}$
@@
$\leqslant a_{0}\rho(x,y)^{v_{0}}/(\lambda^{v_{0}}-1)$
for each $\overline{x}\in T^{-n}(x)$. Together with (2.6), we can conclude that
$\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)$ $=\sum_{\overline{x}\in T^{-n}(x)}\exp(S_{n}\overline{\varphi}(\overline{x}))$
$\leqslant\sum_{\overline{x}\in T^{-n}(x)}\exp(S_{n}\overline{\varphi}(y(\overline{x})))\cdot\exp(a_{0}\rho(x,y)^{v_{0}}/(\lambda^{v_{0}}-1))$
$\leqslant\exp(a_{0}\rho(x,y)^{v_{0}}/(\lambda^{v_{0}}-1))\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(y)$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. ∎
@@
Fix arbitrary constants $\eta,\,\xi,\,a_{0},\,v_{0}>0$, and $\lambda>1$. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2 with constants $\eta$, $\lambda$, and $\xi$, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ be Hölder continuous with respect to the metric $\rho$ with exponent $v_{0}$ and constant $a_{0}$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$, $a\coloneqq\frac{a_{0}}{\lambda^{v_{0}}-1}$, and $D\coloneqq\max_{x\in X}\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}$. Denote by $G=\{x_{1},\,x_{2},\,\ldots,\,x_{l}\}$ a $\xi$-net of $X$. Then for all $x,\,y\in X$, and $n\in\mathbb{N}$, we have
(2.10) $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\leqslant C\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)$
and
(2.11) $C^{-1}\leqslant\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(x)\leqslant C.$
Here $C\coloneqq D^{N}\exp(4a\xi^{v_{0}}+2N\|\varphi\|_{\infty})>1$, where $N\in\mathbb{N}$ is a constant satisfying that
@@
See [PU10, Theorem 4.3.12 (2)] for the existence of $N$.
@@
(2.13) $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\big{/}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)\leqslant\exp(a\rho(x,y)^{v_{0}})\leqslant\exp(a\xi^{v_{0}}).$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$ and each $n\in\mathbb{N}$. Now we are ready to demonstrate (2.10). Let $x,\,y\in X$. Since $G$ is a $\xi$-net, there exist two integers $i,\,j\in[1,l]$ satisfying that $\rho(x_{i},x)<\xi$ and $\rho(x_{j},y)<\xi$. Hence, by (2.12), there exists an integer $m\in[0,N]$ such that $T^{m}(B(x_{i},\xi))\cap B(x_{j},\xi)\neq\emptyset$. Now let $u$ be a point in $T^{-m}(B(x_{j},\xi))\cap B(x_{i},\xi)$. By (2.6) and (2.13), then for each $n\in\mathbb{N}$, we have
@@
$=\exp(a\xi^{v_{0}})\sum_{\overline{u}\in T^{-n}(u)}\exp(S_{n+m}\overline{\varphi}(\overline{u})-S_{m}\overline{\varphi}(u))$
$\leqslant\exp\bigl{(}a\xi^{v_{0}}-m\inf_{x\in X}\overline{\varphi}(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+m)}(T^{m}(u))}\exp(S_{n+m}\overline{\varphi}(\overline{u}))$
$=\exp\bigl{(}a\xi^{v_{0}}-m\inf_{x\in X}\overline{\varphi}(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+m)}(T^{m}(u))}\exp(S_{n}\overline{\varphi}(T^{m}(\overline{u}))+S_{m}\overline{\varphi}(\overline{u}))$
$\leqslant\exp\bigl{(}a\xi^{v_{0}}+m\sup_{x\in X}\overline{\varphi}(x)-m\inf_{x\in X}\overline{\varphi}(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+m)}(T^{m}(u))}\exp(S_{n}\overline{\varphi}(T^{m}(\overline{u})))$
$\leqslant\exp(a\xi^{v_{0}}+2N\|\overline{\varphi}\|_{\infty})D^{N}\sum_{\overline{u}\in T^{-n}(T^{m}(u))}\exp(S_{n}\overline{\varphi}(\overline{u}))$
$=\exp(a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(T^{m}(u))$
@@
Therefore, by (2.13), it follows from $\rho(x_{i},x)<\xi$ and $\rho(x_{j},y)<\xi$ that for each $n\in\mathbb{N}$,
$\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)$ $\leqslant\exp(a\xi^{v_{0}})\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x_{i})$
$\leqslant\exp(3a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x_{j})$
$\leqslant\exp(4a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y).\qed$
By (2.7), for each $n\in\mathbb{N}$,
@@
This implies that $\inf_{x\in X}\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(x)\leqslant 1\leqslant\sup_{x\in X}\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(x).$ Then by (2.10), we obtain (2.11) for all $n\in\mathbb{N}$ and $x\in X$.
@@
Fix arbitrary constants $\eta,\,\xi,\,a_{0},\,v_{0}>0$, and $\lambda>1$. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2 with constants $\eta$, $\lambda$, and $\xi$, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ be Hölder continuous with respect to the metric $\rho$ with exponent $v_{0}$ and constant $a_{0}$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$. Then the sequence $\bigl{\{}\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})\bigr{\}}_{n\in\mathbb{N}}$ converges uniformly to a function $u_{\varphi}\in C(X)$ satisfying
@@
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$, and
@@
To prove this corollary, we first demonstrate (2.15), (2.16), (2.17), and (2.18) for a subsequential limit of the sequence $\bigl{\{}\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})\bigr{\}}_{n\in\mathbb{N}}$, then prove that this sequence has a unique subsequential limit, and finally establish that $\mu$ is the unique $T$-invariant Gibbs state for $T$ and $\varphi$. First, by (2.9) and (2.11), we have

### v3

@@ 2.2. Basic concepts in ergodic theory
where $\mathcal{V}$ ranges over all covers of $X$ that are refinements of the cover $(\mathcal{V}_{n})_{T}^{m}$. In particular, if the potential $\varphi$ is identically zero, then the topological pressure $P(T,\varphi)$ is also called the topological entropy of $T$, and is denoted by $h_{\operatorname{top}}(T)$.
Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2, and $\mu\in\mathcal{M}(X,T)$ be a $T$-invariant Borel probability measure. Now we define measure-theoretic entropy $h_{\mu}(T)$ of $T$ with respect to $\mu$. Recall that a $\mu$-measurable function $J\colon X\rightarrow[0,+\infty)$ is a Jacobian of $T$ with respect to $\mu$ if for each $\mu$-measurable set $A\subseteq X$ on which $T$ is injective, $T(A)$ is still $\mu$-measurable and $\mu(T(A))=\int_{A}\!J\,\mathrm{d}\mu.$ By [PU10, Proposition 2.9.5], there exists a unique Jacobian $J_{\mu}$ of $T$ with respect to $\mu$. Then the measure-theoretic entropy of $T$ with respect to $\mu$ can be defined by $h_{\mu}(T)\coloneqq\int\!\log\bigl{(}J_{\mu}\bigr{)}\,\mathrm{d}\mu$ (see e.g. [PU10, Proposition 2.9.7]).
Under the above notations, we have the following result (see e.g. [PU10, Theorem 3.4.1]).
@@
Let $(X,\rho)$ be a compact metric space. Then for each continuous map $T\colon X\rightarrow X$ and each continuous function $\varphi\colon X\rightarrow\mathbb{R}$, we have
$P(T,\varphi)=\sup\left\{h_{\mu}(T)+\int\!\varphi\,\mathrm{d}\mu:\mu\in\mathcal{M}(X,T)\right\}.$
In particular, if $\varphi$ is identically zero, then $h_{\operatorname{top}}(T)=\sup\{h_{\mu}(T):\mu\in\mathcal{M}(X,T)\}$.
@@
Let $(X,\rho)$ be a compact metric space. Then for each continuous map $T\colon X\rightarrow X$ and each continuous function $\varphi\colon X\rightarrow\mathbb{R}$, a measure $\mu\in\mathcal{M}(X,T)$ is called an equilibrium state for $T$ and $\varphi$ if
@@
We denote by $\mathcal{E}(T,\varphi)$ the set of all equilibrium states for $T$ and $\varphi$. In particular, if $\varphi$ is identically zero, then an equilibrium state for $T$ and $\varphi$ is also called a measure of maximal entropy of $T$.
@@ 2.3. Ruelle–Perron–Frobenius operators and Gibbs states
In this subsection, we review some definitions and results for Ruelle–Perron–Frobenius operators, Gibbs states, and equilibrium states.
Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2, and $\varphi\colon X\rightarrow\mathbb{R}$ be a continuous function. Recall that the Ruelle operator $\mathcal{L}_{\varphi}$ acting on $C(X)$ is given by
@@
for each $u\in C(X)$ and each $x\in X$.
By [PU10, Theorem 5.2.8], there exists an eigenmeasure $m\in\mathcal{P}(X)$ and an eigenvalue $c>0$ of the adjoint operator $\mathcal{L}_{\varphi}^{*}$ of $\mathcal{L}_{\varphi}$, i.e., $\mathcal{L}^{*}_{\varphi}(m)=cm$. Moreover, by [PU10, Proposition 5.2.11], the constant $c$ is precisely $\exp(P(T,\varphi)).$
It is convenient to consider the normalized Ruelle–Perron–Frobenius operator $\mathcal{L}_{\overline{\varphi}}$, where $\overline{\varphi}$ is given by $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$. Hence, by (2.7), we have
@@
(2.9) $C^{-1}\leqslant\frac{\mu(T^{-n}_{x}(B(T^{n}(x),\xi)))}{\exp(S_{n}\varphi(x)-Pn)}\leqslant C\quad\text{ for all }x\in X\text{ and }n\in\mathbb{N}.$
The following result implies that a $T$-invariant Gibbs state for $T$ and $\varphi$ is an ergodic equilibrium state for $T$ and $\varphi$.
@@
Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2. Then for all $\mu\in\mathcal{M}(X,T)$ and Hölder continuous function $\varphi\colon X\rightarrow\mathbb{R}$, the following statements are equivalent:
@@
    $\mu$ is an ergodic equilibrium state for $T$ and $\varphi$.
@@
[…]
The existence and uniqueness of the subsequential limit $u_{\varphi}$ is well-known. We include a proof in Corollary 2.9 together with some quantitative bounds specialized for the purpose of Computable Analysis.
In the remainder of this subsection, we quantitatively re-develop some familiar facts about the Ruelle–Perron–Frobenius operators from the perspective of Computable Analysis in preparation for more technical analysis in Section 4.
@@
Fix arbitrary constants $\eta,\,\xi,\,a_{0}>0,\,\lambda>1$, and $v_{0}\in(0,1]$. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2 with constants $\eta$, $\lambda$, and $\xi$, and $\varphi\in C^{0,v_{0}}(X,\rho)$ satisfy $a_{0}\geqslant\lvert\varphi\rvert_{v_{0},\rho}$. Then we have
(2.10) $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\leqslant\exp\bigl{(}a\rho(x,y)^{v_{0}}\bigr{)}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)$
for each $n\in\mathbb{N}$ and each pair of $x,\,y\in X$ with $\rho(x,y)<\xi$. Here $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$, and $a\coloneqq\frac{a_{0}}{\lambda^{v_{0}}-1}$.
@@
Fix an integer $n>0$ and a pair of $x,\,y\in X$ with $\rho(x,y)<\xi$. Since $T$ is open and distance-expanding with constants $\eta$, $\lambda$, and $\xi$, for each $\overline{x}\in T^{-n}(x)$, by (2.5), one can find a point $y(\overline{x})\coloneqq T_{\overline{x}}^{-n}(y)\in T^{-n}(y)$ satisfying $\rho\bigl{(}T^{i}(\overline{x}),T^{i}(y(\overline{x}))\bigr{)}\leqslant\lambda^{i-n}\rho(x,y)<\rho(x,y)<\xi$ for each integer $0\leqslant i\leqslant n-1$. Since $\varphi\in C^{0,v_{0}}(X,\rho)$, we have $\overline{\varphi}\in C^{0,v_{0}}(X,\rho)$ with $a_{0}\geqslant\lvert\overline{\varphi}\rvert_{v_{0},\rho}$. Hence, we have
$\left\lvert S_{n}\overline{\varphi}(\overline{x})-S_{n}\overline{\varphi}(y(\overline{x}))\right\rvert$ $\leqslant\sum_{i=0}^{n-1}\left\lvert\overline{\varphi}\bigl{(}T^{i}(\overline{x})\bigr{)}-\overline{\varphi}\bigl{(}T^{i}(y(\overline{x}))\bigr{)}\right\rvert$
$\leqslant\sum_{i=0}^{n-1}a_{0}\rho\bigl{(}T^{i}(\overline{x}),T^{i}(y(\overline{x}))\bigr{)}^{v_{0}}$
@@
$\leqslant a\rho(x,y)^{v_{0}}$
for each $\overline{x}\in T^{-n}(x)$. Together with (2.7), we can conclude
$\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)=\sum_{\overline{x}\in T^{-n}(x)}e^{S_{n}\overline{\varphi}(\overline{x})}\leqslant\sum_{\overline{x}\in T^{-n}(x)}e^{a\rho(x,y)^{v_{0}}}\cdot e^{S_{n}\overline{\varphi}(y(\overline{x}))}\leqslant e^{a\rho(x,y)^{v_{0}}}\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(y).\qed$
@@
Fix arbitrary constants $\eta,\,\xi,\,a_{0}>0,\,\lambda>1$, and $v_{0}\in(0,1]$. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2 with constants $\eta$, $\lambda$, and $\xi$, and $G=\{x_{i}:i\in[1,l]\cap\mathbb{N}\}$ be a $\xi$-net of $X$. Assume that the function $\varphi\in C^{0,v_{0}}(X,\rho)$ satisfies $a_{0}\geqslant\lvert\varphi\rvert_{v_{0},\rho}$. Put $a\coloneqq\frac{a_{0}}{\lambda^{v_{0}}-1}$, $D\coloneqq\max_{x\in X}\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}$, and $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$. Then for all $x,\,y\in X$, and $n\in\mathbb{N}$, we have
(2.11) $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\leqslant C\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)\qquad\text{ and }$
(2.12) $C^{-1}\leqslant\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(x)\leqslant C,$
where $C\coloneqq D^{N}\exp(4a\xi^{v_{0}}+2N\|\varphi\|_{\infty})>1$. Here $N\in\mathbb{N}$ satisfies that
@@
Since $X$, $\rho$, and $T$ satisfy the Assumption C in Subsection 2.2, there exists an integer $N\in\mathbb{N}$ satisfying (2.13) and $D<+\infty$.
@@
(2.14) $\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\big{/}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y)\leqslant\exp(a\rho(x,y)^{v_{0}})\leqslant\exp(a\xi^{v_{0}})$
for each pair of $x,\,y\in X$ with $\rho(x,y)<\xi$ and each $n\in\mathbb{N}$.
First, we establish (2.11). To accomplish this, we fix an integer $n$ and a pair of $x,\,y\in X$. Since $G$ is a $\xi$-net, there exist two integers $i,\,j\in[1,l]$ with $\rho(x_{i},x)<\xi$ and $\rho(x_{j},y)<\xi$. Hence, by (2.13), there exists an integer $N_{0}\in[0,N]$ with $T^{N_{0}}(B(x_{i},\xi))\cap B(x_{j},\xi)\neq\emptyset$. Let $u\in T^{-N_{0}}(B(x_{j},\xi))\cap B(x_{i},\xi)$. Then by (2.7) and (2.14), we have
@@
$=\exp(a\xi^{v_{0}})\sum_{\overline{u}\in T^{-n}(u)}\exp(S_{n+N_{0}}\overline{\varphi}(\overline{u})-S_{N_{0}}\overline{\varphi}(u))$
$\leqslant\exp\bigl{(}a\xi^{v_{0}}-N_{0}\inf_{x\in X}\overline{\varphi}(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+N_{0})}(T^{N_{0}}(u))}\exp(S_{n+N_{0}}\overline{\varphi}(\overline{u}))$
$=\exp\bigl{(}a\xi^{v_{0}}-N_{0}\inf_{x\in X}\overline{\varphi}(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+N_{0})}(T^{N_{0}}(u))}\exp\bigl{(}S_{n}\overline{\varphi}\bigl{(}T^{N_{0}}(\overline{u})\bigr{)}+S_{N_{0}}\overline{\varphi}(\overline{u})\bigr{)}$
$\leqslant\exp\bigl{(}a\xi^{v_{0}}+N_{0}\sup_{x\in X}\overline{\varphi}(x)-N_{0}\inf_{x\in X}\overline{\varphi}(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+N_{0})}(T^{N_{0}}(u))}\exp\bigl{(}S_{n}\overline{\varphi}\bigl{(}T^{N_{0}}(\overline{u})\bigr{)}\bigr{)}$
$=\exp\bigl{(}a\xi^{v_{0}}+N_{0}\sup_{x\in X}\varphi(x)-N_{0}\inf_{x\in X}\varphi(x)\bigr{)}\sum_{\overline{u}\in T^{-(n+N_{0})}(T^{N_{0}}(u))}\exp\bigl{(}S_{n}\overline{\varphi}\bigl{(}T^{N_{0}}(\overline{u})\bigr{)}\bigr{)}$
$\leqslant\exp(a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\sum_{\overline{u}\in T^{-n}(T^{N_{0}}(u))}\exp(S_{n}\overline{\varphi}(\overline{u}))$
$=\exp(a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})\bigl{(}T^{N_{0}}(u)\bigr{)}$
@@
Hence, combining (2.14) with $\rho(x_{i},x)<\xi$ and $\rho(x_{j},y)<\xi$, it follows that for each $n\in\mathbb{N}$,
$\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x)\leqslant\exp(a\xi^{v_{0}})\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x_{i})$ $\leqslant\exp(3a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(x_{j})$
$\leqslant\exp(4a\xi^{v_{0}}+2N\|\varphi\|_{\infty})D^{N}\mathcal{L}_{\overline{\varphi}}^{n}(\mathbbm{1})(y).$
Thus, we complete the proof of (2.11). By (2.8), for each $n\in\mathbb{N}$, we have
@@
Then by (2.15), we have that $\inf_{x\in X}\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(x)\leqslant 1\leqslant\sup_{x\in X}\mathcal{L}^{n}_{\overline{\varphi}}(\mathbbm{1})(x).$ Hence, by (2.11), we obtain (2.12) for all $n\in\mathbb{N}$ and $x\in X$. ∎
@@
Fix arbitrary constants $\eta,\,\xi,\,a_{0}>0,\,\lambda>1$, and $v_{0}\in(0,1]$. Let $X,\,\rho,\,T$ satisfy the Assumption C in Subsection 2.2 with constants $\eta$, $\lambda$, and $\xi$, and $\varphi\in C^{0,v_{0}}(X,\rho)$ satisfy $a_{0}\geqslant\lvert\varphi\rvert_{v_{0},\rho}$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$. Then the sequence $\bigl{\{}\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})\bigr{\}}_{n\in\mathbb{N}}$ converges uniformly to a function $u_{\varphi}\in C(X)$ satisfying
@@
for each pair of $x,\,y\in X$ with $\rho(x,y)<\xi$, and
@@
To prove this corollary, we first demonstrate (2.16), (2.17), (2.18), and (2.19) for a subsequential limit of the sequence $\bigl{\{}\frac{1}{n}\sum_{j=0}^{n-1}\mathcal{L}_{\overline{\varphi}}^{j}(\mathbbm{1})\bigr{\}}_{n\in\mathbb{N}}$, then prove that this sequence has a unique subsequential limit. Finally, we show that $\mu$ is the unique $T$-invariant Gibbs state for $T$ and $\varphi$.
First, by (2.10) and (2.12), we have